OpenStudy (anonymous):

4 questions about SQUARE ROOTS.... Can anyone help me!?

OpenStudy (anonymous):

1. Simplify the Square Root (it's a: - sign in front of the S.R. symbol –0.5 5 –5 –25

OpenStudy (anonymous):

\[-\sqrt{25}\] you mean this?

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

-5

jhonyy9 (jhonyy9):

-5

OpenStudy (anonymous):

2. Estimate the square root to the nearest integer. Find two integers that make the equation to make it [image] true. 5, –5 5, 0.5 \[-\sqrt{5}\], \[\sqrt{5}\] 5, 25

OpenStudy (anonymous):

5, –5

jhonyy9 (jhonyy9):

y_1,2=+/- 5

OpenStudy (anonymous):

Which of these sets of numbers contains no irrational numbers?

OpenStudy (anonymous):

Sorry.. wrong attachment

OpenStudy (anonymous):

1º one

OpenStudy (anonymous):

Which number is irrational

OpenStudy (anonymous):

sqrt156

OpenStudy (anonymous):

Okay thanks sooo much!!!!